OpenStudy (anonymous):

If √6 is the geometric mean between 6 and another number, then the number is 1 2 6

OpenStudy (anonymous):

if two numbers are x and y. Then geometric mean = sqrt (x*y) let the other number be x, hence\[\sqrt{6x} = \sqrt{6}\] Hence x = 1